OpenStudy (anonymous):

If a string is 38 millimeters long, how many meters long is it?

Parth (parthkohli):

1 millimeter = 1/10 cm. 1 cm = 1/100 m. 1 millimeter = (1/10 * 1/100) m

Parth (parthkohli):

38 millimeters = 38 * (1/10 * 1/100) meters

OpenStudy (shubhamsrg):

woah! :O

Parth (parthkohli):

yeh! :O

OpenStudy (anonymous):

it is 0.038 m

OpenStudy (anonymous):

thanks

OpenStudy (shubhamsrg):

welcome! ^_^